A sniper holding a Barret M82 Sniper Rifle fires a bullet with a speed of 853 m/s at an
angle of 35° above the horizontal.

a. What height does the bullet reach?
b. How long is the bullet in the air?
c. What is the horizontal range?
1
Expert's answer
2020-10-05T15:30:56-0400

As per the given question,

Speed of the bullet "(u)=853 m\/sec"

Angle "(\theta)=35^\circ"

a) Height reach by the bullet "(H_{max})=\frac{u^2\sin^2\theta}{2g}"


Now substituting the values, "(H_{max})=\frac{853^2\sin^235^\circ}{2g}"


"\Rightarrow (H_{max})=\frac{727609\times 0.329}{2\times 9.8}m"

"\Rightarrow (H_{max})=12213.43m"


b) Let the bullet stay till T time,

"(t)=\frac{2v\sin\theta}{g}"

Now, substituting the values,

"(t)=\frac{2\times853\times \sin35}{9.8}sec"

"\Rightarrow (t)=99.84sec"


c) Let the horizontal range of the bullet is (R)

"(R)=\frac{v^2\sin2\theta}{g}"

Now, substituting the values,

"(R) =\frac{853^2\times\sin 70}{9.8}m"

"\Rightarrow (R)=69768.24m"